# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

### v1

@@ 3.2 Wave buoy orbital measurements
To determine the buoy orientation along the wave-flume we assume that the buoys movement is mainly varying along the wave direction. Since we do not have a magnetometer we do not expect the $x$ and $y$ axes to be pointing in the same direction on long time-scales (limited by gyroscope drift). Figure 3 shows the vertical motion as a function of time for a sine wave of $0.3$ Hz (Figure 1(d)), as well as the vertical and horizontal acceleration.
In order to obtain the trajectory of the wave buoy we apply principal component analysis (PCA) to the high-pass filtered $x$ and $y$ acceleration. The acceleration is filtered to remove drift and only consider the movement due to wave motion. The first principal component, $\bf u_{0}$, is the direction in the $xy$-plane along which the variation is greatest, the second component ($\bf u_{1}$) is orthogonal to $\bf u_{0}$ representing the rest of the variance.
We project the $x$ and $y$-accelerations onto the normalized first principal component ($\frac{\bf u_{0}}{\left\|\bf u_{0}\right\|}$, horizontal) and the normalized second principal component ($\frac{\bf u_{1}}{\left\|\bf u_{1}\right\|}$, vertical), the original movement in the $xy$-plane and the re-projected movement is shown in Figure 4:
@@
If the buoy is only moving along one direction, and the gyroscope is not drifting, all movement should be aligned along the first principal component. It is clearly visible in Figure 4 that this is the case for almost all the movement. The ratio between the two principal components for 10 seconds in the wave-flume is $925$ (upper panel in Figure 4), which provides a measure of how robust this method is on this time scale.
@@
Figure 4: Trajectory of the B25 wave-buoy in the wave-flume with 10 seconds of $0.3$ Hz sinusoidal waves (Figure 1(d)). Upper: xy-plane with principal components and re-projected trajectory. Middle: high-pass filtered trajectory in horizontal–vertical plane along first ($\bf u_{0}$) and second ($\bf u_{1}$) principal component. Lower: full bandwidth trajectory in horizontal–vertical plane for first principal component.
@@
At the surface, $z=0$, Eq. (11) becomes $\tanh(kd)$. The same relation (11) holds true for the acceleration, $\frac{a_{z}}{a_{x}}$ (differentiate Eq. (9) and Eq. (10) twice). In this setup $k=0.71\;\textrm{m}^{-1}$ for $f=0.3$ Hz, and $\tanh(kd)$ is $0.51$, while the orbital motions in Figure 4 (middle panel) have $\frac{u_{z}}{u_{x}}$ of approximately $0.36$.
@@ 3.3 Breaking waves
Two observations of the relation between the measured harmonic sine waves and the breaking waves are of particular interest. First, the maximum horizontal acceleration for the sine wave is of the order of ±$\pm 0.05g$ (Figure 3), whereas for the breaking wave in the lab it exceeds ±$\pm 1g$. Second, buoys 23 and 25 show similar accelerations and vertical displacements for the non-breaking wave (upper panel in Figure 3). However, for a breaking wave the integrated vertical displacement (upper panel in Figure 4) is dominated by exaggerated low-frequency signals for buoy 23. The major difference between the two buoys during the recording of the breaking wave is that buoy 23 is sampling the IMU at $52$ Hz, while buoy 25 is sampling at the full $208$ Hz before filtering and down-samplng to $52$ Hz after feeding and re-projecting the acceleration through the AHRS fusion algorithm. These results will be discussed further in Sec. 66.1.
@@ 4 Open water: Fedjeosen
A wave buoy built into a fender was deployed at $60.752145^{\circ}$ N, $4.693160^{\circ}$ E, moored at $90$ m depth, west of Hellisøy lighthouse in Fedjeosen in July 2023. The location is exposed to to westerly waves from the North Sea, but is sheltered from the north by Fedje Island. The configuration is described below (shown in Figures 6 and 19) and is very similar to recommended moorings described by Martini et al. (2021).
@@
(a) Location of mooring in Fedjeosen, map from Kartverket (sjøkart).
(b) A fender of dimensions $65$ cm x $88$ cm floats at the surface, and houses the electronics and batteries.
Figure 6: Fedjeosen mooring.
@@
Figure 8: Observations from the buoy compared to the simulated $H_{m0}$ (WAM800 with current). Linear regression and quantile-quantile (Q-Q) lines are shown. The model predict slightly higher significant wave height, but with a otherwise fairly linear relationship (Pearson correlation coefficient, $R=0.94$). The model without current shows the same $R$-value.
@@
The measured significant wave height is compared to operational coastal wave models (Fig. 7 and 8), with and without currents (The Wamdi Group, 1988). The nearest grid point in the wave model is about $280$ m away from the buoy location. Figure 8 compares the observations against the simulated $H_{m0}$: there is a good match with an $R$-value of $0.94$ for both models. The buoy measures slightly lower $H_{m0}$ values than model predicts, but the relationship is close to linear.
@@
Deploying the SFY at an exposed location like Fedjeosen provides the opportunity to compare its recordings with satellite observations of significant wave height to add an additional reference for the comparison. Measured $H_{m0}$ was compared to the near real time L3 satellite altimeter measurements from the Copernicus Marine Service catalogue product [^2] WAVE_GLO_PHY_SWH_L3_NRT_014_001 (E.U. Copernicus Marine Service Information . Marine Data Store (MDS)(2023), CMEMS). The satellite data are retrieved and collocated with the SFY using the open source tool wavy [^3] https://wavyopen.readthedocs.org following the collocation routine with temporal and spatial constraints as outlined by Bohlinger et al. (2019). To acquire enough data, we set the temporal constraints to ±$\pm 10$ minutes of the time stamp of the significant wave height, and a maximum spatial distance of $100$ km. This resulted in matches with 6 different satellites, namely Sentinel-3A/B, Cryosat-2, CFOSAT, SARAL-AltiKa, and Sentinel-6A Michael Freilich. Satellite observations appear to follow closely the time series of $H_{m0}$ recorded by the SFY (Fig.7).
(a) Hjeltefjorden ($52$ Hz), logarithmic axes.
(b) Fedjeosen ($208$ Hz), logarithmic axes.
(c) Hjeltefjorden ($52$ Hz), linear axes.
(d) Fedjeosen ($208$ Hz), linear axes.
@@
Figure 9: Elevation energy (variance density) spectra for comparable situations (similar $H_{m0}$ and $T_{p}$) of the same buoy (Figure 5(b)) deployed at two different locations. The IMU (before feeding the AHRS-algorithm) was configured to sample at $52$ Hz in the first case (Hjeltefjorden), and at $208$ Hz in the second case (Fedjeosen). The arrow in the lower left panel shows low-frequency noise caused by insufficient sample rate.
@@
In Figure 9 spectra for 20-minute windows of the time series are compared to spectra for the same buoy deployed inside Hjeltefjorden a few months earlier. The deployment was very similar except that for the first deployment the IMU was sampling at $52$ Hz, while for the second (in Fedjeosen) the IMU was sampling at $208$ Hz and downsampled after being passed through the AHRS-algorithm. The spectra (Figure 9) have been chosen for similar wave situation in terms of $H_{m0}$ and $T_{p}$. A notable difference is that for the buoy configured to sample at the lower frequency ($52$ Hz), more low-frequency noise is observed (arrow in Figure 8(a)). This difference will be discussed more in Sec. 66.1.
@@
Breaking waves hitting or hurling a buoy cause rapid acceleration impulses as can be clearly seen in the wave flume (Figure 5). The strength of this impulse depends on buoy dimensions and wave steepness and how the breaking wave impacts the buoy. Seldal (2023) did a preliminary investigation into the signature of the breaking waves to try distinguish the different types of breakers in a way similar to the approach of Brown and Paasch (2021). The simplest method of quantifying the number of breakers would be to count maxima in the time series of horizontal acceleration, but this would detect many peaks for each breaking event. In order to make the method more robust against false positives we apply the method from Brown et al. (2018). Here we calculate the Short-Time Fourier-Transform (STFT) for tapered windows of $t_{\textrm{bw}}=1.5$ s (overlap between windows of $t_{bw}/2$) for all three components of acceleration ($a_{z}$, $a_{x}$, and $a_{y}$):
@@
Regular, linear waves, do not exist at significant energy levels at frequencies greater than $2$ Hz and we here assume that the energy found at these frequencies is caused by impulsive accelerations from breaking waves. The frequency was chosen so that it is above the consistent band of energy (upper panel in Fig. 9(a)).
(a) Detected breaking events for the moored buoy in Fedjeosen. Upper: The sum of the acceleration spectrum ($A$). Middle: Detected breakers and $\alpha$. Bottom: Breaker fraction ($Q_{b}$) for 20 minute windows.
(b) 2023-10-14T16:00
(c) 2023-10-24T12:00
(d) 2023-11-04T12:15
Figure 10: Breaking events and photos taken towards the buoy position (located to the left behind the hill) at times marked in Figure 9(a). Photos by Norsk Luftambulanse, used with permission.
Breakers are identified by counting peaks of $\alpha$ above a threshold value. The parameters chosen for window length ($t_{\mathrm{bw}}$), $f_{\mathrm{bc}}$ and the threshold for $\alpha$ are dependent on the buoys response to breakers and signal processing (Brown et al., 2018). As in Brown et al. (2018) we use a threshold $\alpha=3.25\,\mathrm{m}\,\mathrm{s}^{-2}$, and rather increase the cut-off frequency to $3.6$ Hz with the argument that only counting high-frequency impulses is a more robust way of detecting the breakers. Along with the plot of $\alpha$ and the detected breakers in the middle panel in Fig. 9(a), we also plot the norm of the acceleration components $\sqrt{a_{z}^{2}+a_{x}^{2}+a_{y}^{2}}$, filtered above $2$ Hz. It gives very similar results to $\alpha$, with the main difference caused by the smoothing due to the windowing in $\alpha$ (middle panel in Fig. 11).
In Figure 11 broadbanded maxima in acceleration (visible as vertical stripes in the acceleration spectrum) are used to detect breakers. In the lower panel the three acceleration components are shown (unfiltered). Similar impulses to the lab experiments are apparent (Fig. 5).
@@
Here, $N_{b}$ is the number of breakers detected during a measurement interval, $\tau=20$ min, and $T$ is the wave period, here approximated with $T_{m01}$. Figure 9(a) shows $Q_{b}$ (lower panel) for a subset of the data. We have visually inspected images taken every 15 minutes with an air traffic camera on the Fedje Island pointing in the direction of the wave buoy. In Figure 10 photos from three instants are shown. These are marked in Figure 9(a) with vertical dashed lines. Qualitatively, the detection of breakers agrees well with the visual impression of breaking waves in the photographs. The wave steepness ($H_{m0}/(g\cdot\frac{T_{m01}^{2}}{2\pi})$) for the measurement period varies very little (but agree well with models).
@@ 5.1 Array of buoys in the surf
(a) Drift trajectory with the instantaneous geometry shown as dashed white lines. Aerial images © norgeibilder.no, Kartverket (Rogaland 2019).
(b) Wave conditions on the day. The buoys are visible on the crest, to the left on the wave that is breaking.
@@
Figure 13: Vertical acceleration spectrograms for three buoys, captured in the surf at Jæren. Vertical lines of high intensity signals spanning the full frequency spectrum is visible at each buoy, with some delay in phase.
Figure 13 shows the acceleration density spectrogram for the vertical component of the three buoys for drift shown in Figure 12. Vertical lines of high energy impulses assumed to be caused by breaking events are visible, and are correlated in time between the buoys. Since the buoys in this experiment were configured with very sensitive accelerometer range (±$\pm 2g$) and gyroscope range (±$\pm 125$ dps), the sensor saturates and the trajectories are difficult to recover.
@@ 5.2 Trajectory in a breaking wave
In a similar experimental setup, a single SFY buoy was deployed in the surf at Green Bowl, Bali, Indonesia, in November 2023. In this deployment the acceleration range is set to ±$\pm 16g$ and the gyroscope range is set to ±$\pm 1000$ dps, and the AHRS filter is run at $208$ Hz. First the spectrum (Figure 14) was measured for 20 minutes outside the breaking waves as a reference. The swell was about 0.9 m while the surfing wave height was about 0.3–0.5 m. The spectrum does not consist of a single peak, so irregular sets of waves breaking onto the shore is expected.
@@
The buoy was then deployed so that it would drift towards the shore through the breaking waves. The horizontal components are re-projected along the direction of maximum variance as in Sec. 33.2 and integrated to the horizontal and vertical displacement. A cut-off frequency of $0.05$ Hz is used when filtering before integrating.
In Figure 16 a wave approaches the buoy (painted red). The buoy is drawn offshore towards the surf wave (a), and almost escapes the breaking crest (b), but eventually the buoy is caught in the overturning wave (c) and thrown towards the surface of the water (exceeding $10g$ upwards acceleration on impact) and flushed out. A second orbit begins as the next (non-breaking) wave travels towards the shore.
@@
(a) 22 seconds
(b) 23 seconds
(c) 24 seconds
(d) Orbital movement of the buoy for the time between vertical lines in Figure 15. The black circle marks the start of the buoy trajectory and each subsequent second is marked by a red circle. The black triangle marks the end.
Figure 16: Orbital movement of buoy and photos captured at the same breaking wave recorded by the buoy in Figure 15.
The recorded displacement amplitudes match well with expected values from the spectrum and visual observations. Usually, measurements from breaking waves in the surf are not suitable to be integrated to displacement without heavy filtering. In Figure 15 we can match the photographs with the integrated trace, and see that the recorded trajectory is sensible even with a very low cut-off frequency. This is discussed further in Sec. 66.2.

### v2

@@ 3.2 Wave buoy orbital measurements
To determine the buoy orientation along the wave-flume we assume that the buoy movement is mainly varying along the wave direction. Since we do not have a magnetometer we do not expect the $x$ and $y$ axes to be pointing in the same direction on long time-scales (limited by gyroscope drift). Figure 3 shows the vertical motion as a function of time for a sine wave of $0.3$ Hz (Fig. 2d), as well as the vertical and horizontal acceleration.
In order to obtain the trajectory of the wave buoy we apply principal component analysis (PCA) to the high-pass filtered $x$ and $y$ acceleration. The acceleration is high-pass filtered (using the Butterworth-filter) so that only movement due to wave motion. The first principal component, $\bf u_{0}$, is the direction in the $xy$-plane along which the variation is greatest, the second component ($\bf u_{1}$) is orthogonal to $\bf u_{0}$ representing the rest of the variance.
We project the $x$ and $y$-accelerations onto the normalized first principal component ($\frac{\bf u_{0}}{\left\|\bf u_{0}\right\|}$, horizontal) and the normalized second principal component ($\frac{\bf u_{1}}{\left\|\bf u_{1}\right\|}$, vertical), the original movement in the $xy$-plane and the re-projected movement is shown in Fig. 4:
@@
If the buoy is only moving along one direction, and the gyroscope is not drifting, all movement should be aligned along the first principal component. It is clearly visible in Fig. 4 that this is the case for almost all the movement. The ratio between the two principal components for 10 seconds in the wave-flume is $925$ (upper panel in Fig. 4), which provides a measure of how robust this method is on this time scale.
@@
Figure 4: Trajectory of the B25 wave-buoy in the wave-flume with about 3 cycles of a $0.3$ Hz (10 seconds) sinusoidal wave (Fig. 2d). $f_{0}$ is the lower cut-off frequency in the bandpass filter when integrating the acceleration and computing the principal components. Upper: xy-plane with principal components and re-projected trajectory. Middle: high-pass filtered trajectory in horizontal–vertical plane along first ($\bf u_{0}$) and second ($\bf u_{1}$) principal component. Lower: full bandwidth trajectory in horizontal–vertical plane for first principal component.
@@
At the surface, $z=0$, Eq. (11) becomes $\tanh(kd)$. The same relation (11) holds true for the acceleration, $\frac{a_{z}}{a_{x}}$ (differentiate Eq. (9) and Eq. (10)). In this setup $k=0.71\;\textrm{m}^{-1}$ for $f=0.3$ Hz, and $\tanh(kd)$ is $0.51$, while the orbital motions in Fig. 4 (middle panel) have $\frac{u_{z}}{u_{x}}$ of approximately $0.36$. We note that for $k=0.71$, the wavelength here is no longer short compared to the length of the flume, and perfect match with theory can not be expected since reflections and standing waves may start to become relevant.
@@ 3.3 Breaking waves
Two observations of the relation between the measured harmonic sine waves and the breaking waves are of particular interest. First, the maximum horizontal acceleration for the sine wave is of the order of ±$\pm 0.05g$ (Fig. 3), whereas for the breaking wave in the lab it exceeds ±$\pm 1g$. Second, buoys 23 and 25 show similar accelerations and vertical displacements for the non-breaking wave (upper panel in Fig. 3). However, for a breaking wave the integrated vertical displacement (upper panel in Fig. 5) is dominated by exaggerated low-frequency signals (long period oscillations) for buoy 23. The major difference between the two buoys during the recording of the breaking wave is that buoy 23 is sampling the IMU at $52$ Hz, while buoy 25 is sampling at the full $208$ Hz before filtering and downsampling to $52$ Hz after feeding and re-projecting the acceleration through the AHRS fusion algorithm. These results will be discussed further in Sec. 66.1.
@@ 4 Open water: Fedjeosen
A wave buoy built into a fender was deployed at $60.752145^{\circ}$ N, $4.693160^{\circ}$ E (Figure 19), moored at $90$ m depth, west of Hellisøy lighthouse in Fedjeosen in July 2023. The location is exposed to westerly waves from the North Sea, but is sheltered from the north by Fedje Island. The configuration is described below (shown in Figures 6 and 20) and is very similar to recommended moorings described by Martini et al. (2021).
@@
Figure 6: Fedjeosen mooring: A fender of dimensions $65$ cm x $88$ cm floats at the surface, and houses the electronics and batteries.
@@
Figure 8: Observations from the buoy compared to the simulated $H_{m0}$ (WAM800 with current). Linear regression and quantile-quantile (Q-Q) lines are shown. The model predict slightly higher significant wave height, but with a otherwise fairly linear relationship (Pearson correlation coefficient, $R=0.96$). The model without current shows the same $R$-value.
We use output from the operational coastal wave model WAM (Günther et al., 1992; The Wamdi Group, 1988; Komen et al., 1996) as a reference for the wave conditions at the buoy site. The wave model WAM (cycle 4.7) is run on a rotated grid with 800m grid spacing for five domains to cover the Norwegian coast. The model is run with ST4 physics (Ardhuin et al., 2010) with a spectral resolution of 36 frequencies and 36 directions forced with surface winds from Arome (Müller et al., 2017). The domains are nested into a WaveWatchIII [^3] https://github.com/NOAA-EMC/WW3.git setup on a large domain with 4 km grid spacing. A separate set of forecasts with WAM are run with input of ocean surface (0 m) currents from the ocean model ROMS [^4] https://github.com/myroms/roms.git. The archived forecasts from Arome, WaveWatchIII, ROMS and WAM can be accessed from an OPeNDAP server [^5] https://thredds.met.no/thredds/fou-hi/fou-hi.html. The first 12 hours of each WAM model run (00 and 12 UTC) at the nearest grid point is used to compile the time series of significant wave height.
@@
The measured significant wave height is compared to the output of significant wave height from the WAM domain for the west coast of Norway, Vestlandet (Fig. 7 and 8), with and without currents. The nearest grid point in the wave model is about $280$ m away from the buoy location. Figure 8 compares the observations against the simulated $H_{m0}$, and there is a good match with an $R$-value of $0.96$ for both models. The buoy measures slightly lower $H_{m0}$ values than model predicts, except for a few instances of higher wave heights where the buoy measures higher than the model predicts. In general the relationship is close to linear.
@@
Deploying the SFY at an exposed location like Fedjeosen provides the opportunity to compare its recordings with satellite observations of significant wave height to add an additional reference for the comparison. Measured $H_{m0}$ was compared to the near real time L3 satellite altimeter measurements from the Copernicus Marine Service catalogue product [^6] https://data.marine.copernicus.eu/product/WAVE_GLO_PHY_SWH_L3_NRT_014_001/description (E.U. Copernicus Marine Service Information , Marine Data Store (MDS)(2023), CMEMS). The satellite data are retrieved and collocated with the SFY using the open source tool wavy [^7] https://wavyopen.readthedocs.org following the collocation routine with temporal and spatial constraints as outlined by Bohlinger et al. (2019). To acquire enough data, we set the temporal constraints to ±$\pm 10$ minutes of the time stamp of the significant wave height, and a maximum spatial distance of $100$ km. This resulted in matches with 6 different satellites, namely Sentinel-3A/B, Cryosat-2, CFOSAT, SARAL-AltiKa, and Sentinel-6A Michael Freilich. Satellite observations appear to follow closely the time series of $H_{m0}$ recorded by the SFY (Fig.7).
@@
Figure 9: Elevation energy (variance density) spectra for comparable situations (similar $H_{m0}$ and $T_{p}$) of the same buoy (Fig. 6) deployed at two different locations. The IMU (before feeding the AHRS-algorithm) was configured to sample at $52$ Hz in the first case (Hjeltefjorden), and at $208$ Hz in the second case (Fedjeosen). The arrow in the lower left panel shows low-frequency noise caused by insufficient sample rate.
@@
In Fig. 9 spectra for 20-minute windows of the time series are compared to spectra for the same buoy deployed inside Hjeltefjorden a few months earlier. The deployment was very similar except that for the first deployment the IMU was sampling at $52$ Hz, while for the second (in Fedjeosen) the IMU was sampling at $208$ Hz and downsampled after being passed through the AHRS-algorithm. The spectra (Fig. 9) have been chosen for similar wave situation in terms of $H_{m0}$ and $T_{p}$. A notable difference is that for the buoy configured to sample at the lower frequency ($52$ Hz), more low-frequency noise is observed (arrow in Fig. 9c). This difference will be discussed more in Sec. 66.1.
@@
Breaking waves hitting or hurling a buoy cause rapid acceleration impulses as can be clearly seen in the wave flume (Fig. 5). A breaking wave strikes the buoy or flushes it sideways, causing an impulsive broad-banded signal that is not dependent on the geometric cut-off frequency. The measured strength of this impulse depends on buoy dimensions and wave steepness and how the breaking wave impacts the buoy, rather than the wave-length of the wave (before it breaks). Seldal (2023) did a preliminary investigation into the signature of the breaking waves to try distinguish the different types of breakers in a way similar to the approach of Brown and Paasch (2021). The simplest method of quantifying the number of breakers would be to count maxima in the time series of horizontal acceleration, but this would detect many peaks for each breaking event. In order to make the method more robust against false positives we apply the method from Brown et al. (2018). Here we calculate the Short-Time Fourier-Transform (STFT) for tapered windows of $t_{\textrm{bw}}=1.5$ s (overlap between windows of $t_{bw}/2$) for all three components of acceleration ($a_{z}$, $a_{x}$, and $a_{y}$):
@@
Regular, linear waves, do not exist at significant energy levels at frequencies greater than $2$ Hz and we here assume that the energy found at these frequencies is caused by impulsive accelerations from breaking waves. The frequency was chosen so that it is above the consistent band of energy (upper panel in Fig. 10).
Figure 10: Breaking events and photos taken towards the buoy position (located to the left behind the hill) at times marked in (a). Photos by Norsk Luftambulanse, used with permission.
Breakers are identified by counting peaks of $\alpha$ above a threshold value. The parameters chosen for window length ($t_{\mathrm{bw}}$), $f_{\mathrm{bc}}$ and the threshold for $\alpha$ are dependent on the buoys’ response to breakers, and signal processing (Brown et al., 2018). As in Brown et al. (2018) we use a threshold $\alpha=3.25\,\mathrm{m}\,\mathrm{s}^{-2}$, and rather increase the cut-off frequency to $3.6$ Hz with the argument that only counting high-frequency impulses is a more robust way of detecting the breakers. Along with the plot of $\alpha$ and the detected breakers in the middle panel in Fig. 10, we also plot the norm of the acceleration components $\sqrt{a_{z}^{2}+a_{x}^{2}+a_{y}^{2}}$, filtered above $2$ Hz. It gives very similar results to $\alpha$, with the main difference caused by the smoothing due to the windowing in $\alpha$ (middle panel in Fig. 11).
In Fig. 11 broadbanded maxima in acceleration (visible as vertical stripes in the acceleration spectrum) are used to detect breakers. In the lower panel the three acceleration components are shown (unfiltered). Similar impulses to the lab experiments are apparent (Fig. 5).
@@
Here, $N_{b}$ is the number of breakers detected during a measurement interval, $\tau=20$ min, and $T$ is the wave period, here approximated with $T_{m01}$. Figure 10a shows $Q_{b}$ (lower panel) for a subset of the data. We have visually inspected images taken every 15 minutes with an air traffic camera on the Fedje Island pointing in the direction of the wave buoy. In Fig. 10 photos from three instants are shown. These are marked in Fig. 10a with vertical dashed lines. Qualitatively, the detection of breakers agrees well with the visual impression of breaking waves in the photographs. The wave steepness ($H_{m0}/(g\cdot\frac{T_{m01}^{2}}{2\pi})$) for the measurement period varies very little (but agrees well with models).
@@ 5.1 Array of buoys in the surf
Figure 13: Vertical acceleration spectrograms for three buoys, captured in the surf at Jæren. Vertical lines of high intensity signals (marked with arrows) spanning the full frequency spectrum is visible at each buoy, with some delay in phase.
Figure 13 shows the acceleration density spectrogram for the vertical component of the three buoys for drift shown in Fig. 12. Vertical lines of high energy impulses assumed to be caused by breaking events are visible, and are correlated in time between the buoys. Since the buoys in this experiment were configured with very sensitive accelerometer range (±$\pm 2g$) and gyroscope range (±$\pm 125$ dps), the sensor saturates and the trajectories are difficult to recover.
@@ 5.2 Trajectory in a breaking wave
In a similar experimental setup, a single SFY buoy was deployed in the surf at Green Bowl, Bali, Indonesia, in November 2023. In this deployment the acceleration range is set to ±$\pm 16g$ and the gyroscope range is set to ±$\pm 1000$ dps, and the AHRS filter is run at $208$ Hz. First the spectrum (Fig. 14) was measured for 20 minutes outside the breaking waves as a reference. The swell was about 0.9 m while the surfing wave height was about 0.3–0.5 m. The spectrum does not consist of a single peak, so irregular sets of waves breaking onto the shore is expected.
@@
The buoy was then deployed so that it would drift towards the shore through the breaking waves. The horizontal components are re-projected along the direction of maximum variance as in Sec. 33.2 and integrated to the horizontal and vertical displacement. A cut-off frequency of $0.04$ Hz is used when filtering before integrating.
In Fig. 16 a wave approaches the buoy (painted red). The buoy is drawn offshore towards the surf wave (a), and almost escapes the breaking crest (b), but eventually the buoy is caught in the overturning wave (c) and thrown towards the surface of the water (exceeding $10g$ upwards acceleration on impact) and flushed out. A second orbit begins as the next (non-breaking) wave travels towards the shore.
@@
Figure 16: Orbital movement of buoy and photos captured at the same breaking wave recorded by the buoy in Fig. 15.
The recorded displacement amplitudes match well with expected values from the spectrum and visual observations. Usually, measurements from breaking waves in the surf are not suitable to be integrated to displacement without heavy filtering. In Fig. 16 we can match the photographs with the integrated trace (Fig. 15), and see that the recorded trajectory (Fig. 16d) is sensible even with a very low cut-off frequency. This is discussed further in Sec. 66.2.